# Topological Phases of Many-Body Localized Systems: Beyond Eigenstate Order

### v1

@@ D.3.1 Unstable invariants
Any family of unitaries $V:S^{n}\to U(N)$ which commutes with $\rho$ must then decompose as
$V=\bigoplus_{r\text{ irrep of }G}\mathbbm{1}_{d_{r}}\otimes V_{r}.$ (112)
Any deformation of this family which continues to commute with $\rho$ must maintain this structure. Thus, the families $V_{1}({\vec{\theta}})$ and $V_{2}({\vec{\theta}})$ are homotopic through a symmetric homotopy if and only if $V_{1r}({\vec{\theta}})$ and $V_{2r}({\vec{\theta}})$ are homotopic for all $r$. We have a complete set of invariants given by the winding numbers of each $V_{r}$,
$W^{\rho}_{n}[V]=(W_{n}[V_{r}])_{r\text{ irrep of }G}.$ (113)
With some foresight, we choose to identify $W^{\rho}_{n}[V]$ with an element of the module
@@
Then we define the complete invariant
$W^{\rho}_{n}[V]=\sum_{r\text{ irrep of }G}r\cdot W_{n}[V_{r}].$ (115)
QCA are properly projective unitaries, rather than unitaries. $W^{\rho}_{n}[V]$ continues to function as an invariant for $n\neq 1$, as in the case without symmetry, but for $n=1$ we can multiply $V(\theta)$ by $e^{im\theta}$ and change all the winding numbers. Indeed, we find
$W^{\rho}_{1}[e^{im\theta}V]$ $=\sum_{r\text{ irrep of }G}r\cdot W_{1}[e^{im\theta}V_{r}]$ (116)
$=\sum_{r\text{ irrep of }G}r\cdot(mn_{r}+W_{1}[V_{r}])$ (117)
$=\left(\sum_{r\text{ irrep of }G}r\otimes\mathbbm{1}_{n_{r}}\right)\cdot m+W^{\rho}_{1}[V]$ (118)
$=\rho\cdot m+W^{\rho}_{1}[V],$ (119)
@@
for $n_{r}\geq 0$ and the decomposition of $\rho$ into irreps, $\rho=\sum_{r}r\otimes\mathbbm{1}_{n_{r}}$. Thus, for projective unitaries, we must consider the coset
$w^{\rho}_{1}[V]=W^{\rho}_{1}[V]+\rho\cdot\pi_{1}(U(N))\cong W^{\rho}_{1}[V]+\rho\cdot{\mathbb{Z}}.$ (121)
@@ D.3.2 Stable invariants
With the unstable invariants $W^{\rho}_{n}[V]$ identified, we must now consider the stabilization process of appending ancillae. We have a directed system among symmetric projective unitary groups $PU(N)^{\rho}$ (the subset of $PU(N)$ which commutes with the representation $\rho$) given by
@@
To compute the action of the induced maps $\iota^{*}_{\rho\to\rho\sigma}(W_{n}^{\rho}[V])=W_{n}^{\rho\sigma}[V\otimes\mathbbm{1}]$, we suppose that we have Clebsch-Gordan multiplicities $C_{rr^{\prime}}^{k}\in{\mathbb{Z}}$ such that, for $r,r^{\prime},k$ irreducible,
@@
Then $V\otimes\mathbbm{1}$ decomposes into the irreducible subspaces of $\rho\sigma$ as
@@ D.3.3 Classification
$\pi_{n}(\mathrm{QCA}_{0}^{\mathcal{R}})\cong\left\{\begin{array}[]{l l}\mathcal{K}^{\mathcal{R}}/{\mathbb{Z}}&\quad n=1,\\ 0&\quad n\text{ even,}\\ \mathcal{K}^{\mathcal{R}}&\quad n\text{ odd,}\end{array}\right.$ (143)
where by ${\mathbb{Z}}$ we mean the subgroup generated by $1\in\mathcal{K}^{\mathcal{R}}$.
@@
Let us suppose that the monoid of ancillae $\mathcal{R}$ contains all finite dimensional representations of $U(1)$, so $\mathcal{R}$ is the full representation rig and thus isomorphic to Laurent polynomials with non-negative coefficients,
@@ D.4.2 Stable invariants
The invariants of the previous section are unstable to adding ancillae, and apparently depend on position $x$. Stabilizing the invariants through the algebraic localization construction fixes both of these issues. We modify the $\tilde{v}_{x}X_{G}$ invariant to
@@ D.4.3 Classification
We have a complete set of invariants in $\mathcal{R}^{-1}\mathrm{Rep}(G)/X_{G}$. Again, not all $\tilde{v}/\rho$ can actually be achieved with a fixed $\mathcal{R}$. The only numerators $\tilde{v}$ which may occur are those which occur as tensor factors in $\rho$ for some $\rho\in\mathcal{R}$. That is, the numerators belong to the saturation
@@
The full classifying group is then
$\pi_{0}(\mathrm{QCA}_{1}^{\mathcal{R}})\cong H^{2}(G;U(1))\oplus\log\mathcal{R}^{-1}\hat{\mathcal{R}}/X_{G}.$ (171)
We express the second factor with a logarithm just to make the group operation addition. The induced homomorphism from evaluation on a product state, $e_{0}^{*}$, just corresponds to projecting to the first summand, and thus reproduces the group cohomology classification of invertible states in one dimension [66, 93].
@@ E.1 Modified swindle
(where $u_{j,j+1}$ is defined in Eq. (7)) for $t\in[0,\tfrac{1}{3})$,
@@ E.2.1 Invariant
The invariant associated to $v(\theta)$ is a coset (Eq. (135)) of formal ratios of representations with the numerator weighted by the winding numbers. We identify the charge 0 representation with 1 and the charge 1 representation with $x\in{\mathbb{N}}[x,x^{-1}]\cong\mathrm{Rep}(U(1))$. Then the invariant is
@@
Indeed, we believe the odd $W$ phase without symmetry to be the two-tone-driven phase of Refs. [29, 30]. While the solvable limit of the model in Ref. [29] is highly singular, which makes computing invariants from integral formulae inconvenient, the edge phenomenology described in Ref. [29] is consistent with what we expect from having $\tilde{w}_{1}\neq 0\bmod 1$. Namely, the edge dynamics was argued to be anomalous in that it could not be achieved by a zero-dimensional localized (in the sense that a continuous micromotion operator exists as in Eq. (35)) system which did not occur at the edge of an extended one-dimensional system. As the model in Ref. [29] is based around a version of the AKLT model, which can be trivialized when stacked with itself, it is thus reasonable to suppose that $\tilde{w}_{1}=1/2\bmod 1$ for that model, as this is the only element of ${\mathbb{Q}}/{\mathbb{Z}}$ of order 2. Thus, the abstract classification reveals a relationship between the interacting QP pump and Refs. [29, 30] which was not clear previously. It also extends the arguments for stability in Ref. [29] to the nonperturbative regime, where a picture of emergent symmetry (as employed in that reference) becomes inapplicable.

### v2

@@ D.3.1 Unstable invariants
We first describe the classification of symmetric unitaries, before explaining how this classification is modified for projective unitaries. Any family of unitaries $\tilde{V}:S^{n}\to U(N)$ which commutes with $\rho$ must decompose as
$\tilde{V}=\bigoplus_{r\text{ irrep of }G}\mathbbm{1}_{d_{r}}\otimes\tilde{V}_{r}.$ (112)
Any deformation of this family which continues to commute with $\rho$ must maintain this structure. Thus, the families $\tilde{V}_{1}({\vec{\theta}})$ and $\tilde{V}_{2}({\vec{\theta}})$ are homotopic through a symmetric homotopy if and only if $\tilde{V}_{1r}({\vec{\theta}})$ and $\tilde{V}_{2r}({\vec{\theta}})$ are homotopic for all $r$. We have a complete set of invariants given by the winding numbers of each $\tilde{V}_{r}$,
$W^{\rho}_{n}[\tilde{V}]=(W_{n}[\tilde{V}_{r}])_{r\text{ irrep of }G}.$ (113)
With some foresight, we choose to identify $W^{\rho}_{n}[\tilde{V}]$ with an element of the module
@@
Then we define the complete invariant for symmetric unitaries
$W^{\rho}_{n}[\tilde{V}]=\sum_{r\text{ irrep of }G}r\cdot W_{n}[\tilde{V}_{r}].$ (115)
QCA are properly projective unitaries, rather than unitaries. $W^{\rho}_{n}[\tilde{V}]$ continues to function as a complete invariant for $n\geq 2$, as in the case without symmetry. We may thus denote them $W^{\rho}_{n}[V]$, where $V:S^{n}\to PU(N)$ is a family of symmetric projective unitaries which lie in the connected component of the identity. However, there are symmetric projective unitaries $V$ which cannot be expressed as symmetric proper unitaries (requiring a larger group of $n=0$ invariants), and there are loops $V(\theta)$ of symmetric proper unitaries which become equivalent when regarded as symmetric projective unitaries (requiring a smaller group of $n=1$ invariants).
We first recall that connected components of symmetric projective unitaries [$n=0$] are labelled by one-dimensional representations of $G$. Given a symmetric projective unitary $V$, we can arbitrarily choose a proper unitary $\tilde{V}$ in the preimage of $V$ under the quotient by global phases. The fact that $V$ is symmetric only implies that $\tilde{V}$ is symmetric up to a global phase,
$\rho(g)^{\dagger}\tilde{V}\rho(g)=\chi[V](g)\tilde{V}.$ (116)
Note that $\chi[V]$ is the same when $\tilde{V}$ is multiplied by a global phase, justifying the notation $\chi[V]$ rather than $\chi[\tilde{V}]$. Using that $\rho$ is a representation, we see that $\chi[V]:G\to U(1)$ is a one-dimensional representation of $G$. From Eq. (110), we see that $\tilde{V}$ must decompose as
$\tilde{V}=\sum_{r\text{ irrep of }G}\sigma_{r\to\chi[V]r}\otimes\tilde{V}_{r},$ (117)
where $\sigma_{r\to\chi[V]r}$ is a matrix with a single nonzero element equal to 1, such that $\tilde{V}$ maps the irrep $r$ block of $\rho$ to the irrep $\chi[V]r$ block. Any path of projective unitaries $V(t)$ lifts to a path of proper unitaries $\tilde{V}(t)$ (as the unit interval is contractible), and the space of permutations $\sigma_{r\to\chi[V]r}$ is discrete, so the permutation label $\chi[V]$ is an invariant for path components of projective unitaries. Any unitaries with the same $\chi[V]$ in the decomposition Eq. (117) can be connected by a path, simply by connecting each $\tilde{V}_{r}$ by a path. Thus, $\chi[V]$ is a complete invariant.
However, not all $\chi[V]\in\mathrm{Hom}(G\to U(1))=:X_{G}$ can occur in a finite Hilbert space. For the $\tilde{V}$ in Eq. (117) to be unitary, we need that $\sum_{r}\sigma_{r\to\chi[V]r}$ is a finite-dimensional permutation matrix, and so in particular $\chi[V]$ has finite order: there is an $n$ such that $\chi[V](g)^{n}=1$ for all $g$. Thus, $\chi[V]$ must belong to the torsion subgroup of $X_{G}$ (the subgroup of elements of finite order, with the group operation in $X_{G}$ defined pointwise), which we denote $X_{G}^{T}$. Conversely, given any $\chi[V]\in X_{G}^{T}$ with order $n$, we can take
$\rho=\bigoplus_{p=0}^{n-1}\chi[V]^{p}$ (118)
and then the unitary $\sum_{p=0}^{n-1}\sigma_{\chi[V]^{p}\to\chi[V]^{p+1}}$ achieves the invariant $\chi[V]$.
As an aside related to the discussion in Appendix D.2, we note that
$X_{G}^{T}\cong H^{1}(G;{\mathbb{Q}}/{\mathbb{Z}})\cong H^{1}(G;\pi_{1}(\mathrm{QCA}_{0}))$ (119)
for the vast majority of groups $G$ relevant to physics. For instance, a sufficient condition is that the component group $G/G^{0}$ (where $G^{0}$ is the identity component of the topological group $G$) is finitely generated.
For $n=1$, we can multiply a loop of proper unitaries $\tilde{V}(\theta)$ by $e^{im\theta}$ and change all the winding numbers, while still forming the same loop of projective unitaries after the quotient by global phases. Indeed, we find
$W^{\rho}_{1}[e^{im\theta}\tilde{V}]$ $=\sum_{r\text{ irrep of }G}r\cdot W_{1}[e^{im\theta}\tilde{V}_{r}]$ (120)
$=\sum_{r\text{ irrep of }G}r\cdot(mn_{r}+W_{1}[\tilde{V}_{r}])$ (121)
$=\left(\sum_{r\text{ irrep of }G}r\otimes\mathbbm{1}_{n_{r}}\right)\cdot m+W^{\rho}_{1}[\tilde{V}]$ (122)
$=\rho\cdot m+W^{\rho}_{1}[\tilde{V}],$ (123)
@@
for $n_{r}\geq 0$ and the decomposition of $\rho$ into irreps, $\rho=\sum_{r}r\otimes\mathbbm{1}_{n_{r}}$. Thus, we must consider the coset
$w^{\rho}_{1}[V]=W^{\rho}_{1}[\tilde{V}]+\rho\cdot\pi_{1}(U(N))\cong W^{\rho}_{1}[\tilde{V}]+\rho\cdot{\mathbb{Z}},$ (125)
which is now a well-defined invariant for loops of projective unitaries $V(\theta)$.
@@ D.3.2 Stable invariants
With the unstable invariants identified, we must now consider the stabilization process of appending ancillae. We have a directed system among symmetric projective unitary groups $PU(N)^{\rho}$ [the subset of $PU(N)$ which commutes with the representation $\rho$] given by
@@
Starting at $n=0$, we find the invariant $\chi[V]=e^{i\omega^{\rho}[V]}$ of $\pi_{0}(PU(N)^{\rho})$ is actually already stable to appending ancillae. With $\omega^{\rho}[V]\in{\mathbb{R}}/(2\pi{\mathbb{Z}})\cong U(1)$ defined modulo $2\pi$, we can see from Eq. (116) that
$\iota_{\rho\to\rho\sigma}^{*}(\omega^{\rho}[V])=\omega^{\rho\sigma}[V\otimes\mathbbm{1}]=\omega^{\rho}[V]$ (128)
and more generally
$\omega^{\rho\sigma}[V\otimes Q]=\omega^{\rho}[V]+\omega^{\sigma}[Q].$ (129)
To compute the action of the induced maps on the remaining unstable invariants, $\iota^{*}_{\rho\to\rho\sigma}(W_{n}^{\rho}[V])=W_{n}^{\rho\sigma}[V\otimes\mathbbm{1}]$, we suppose that we have Clebsch-Gordan multiplicities $C_{rr^{\prime}}^{k}\in{\mathbb{Z}}$ such that, for $r,r^{\prime},k$ irreducible,
@@
Then $V\otimes\mathbbm{1}$ decomposes into the irreducible subspaces of $\rho\sigma$ (possibly after multiplication by a permutation of the irrep subspaces) as
@@ D.3.3 Classification
$\pi_{n}(\mathrm{QCA}_{0}^{\mathcal{R}})\cong\left\{\begin{array}[]{l l}\leq X^{T}_{G}&\quad n=0,\\ \mathcal{K}^{\mathcal{R}}/{\mathbb{Z}}&\quad n=1,\\ 0&\quad n\geq 2\text{ even,}\\ \mathcal{K}^{\mathcal{R}}&\quad n\geq 3\text{ odd,}\end{array}\right.$ (149)
where by ${\mathbb{Z}}$ we mean the subgroup generated by $1\in\mathcal{K}^{\mathcal{R}}$. We have also not precisely characterized the restriction on $\pi_{0}(\mathrm{QCA}_{0}^{\mathcal{R}})\leq X_{G}^{T}$ imposed by the limited ancillae in $\mathcal{R}$, as our focus is on the less familiar higher homotopy groups.
@@
We have that $X_{U(1)}\cong{\mathbb{Z}}$, so $X_{U(1)}^{T}\cong 0$ is trivial. Let us suppose that the monoid of ancillae $\mathcal{R}$ contains all finite dimensional representations of $U(1)$, so $\mathcal{R}$ is the full representation rig and thus isomorphic to Laurent polynomials with non-negative coefficients,
@@ D.4.2 Stable invariants
While the $\omega_{x}$ invariant is stable to stacking ancillae, the $\tilde{v}_{x}X_{G}$ invariant is not, and both apparently depend on position $x$. Stabilizing the $\tilde{v}_{x}X_{G}$ invariant through the algebraic localization construction fixes both of these issues.
We modify the $\tilde{v}_{x}X_{G}$ invariant to
@@ D.4.3 Classification
We have a complete set of invariants in $H^{2}(G;U(1))\times\mathcal{R}^{-1}\mathrm{Rep}(G)/X_{G}$. Again, not all $(\omega,\tilde{v}X_{G}/\rho)$ can actually be achieved with a fixed $\mathcal{R}$. We focus primarily on $\tilde{v}X_{G}/\rho$.
The only numerators $\tilde{v}$ which may occur in the invariant $\tilde{v}X_{G}/\rho$ are those which occur as tensor factors in $\rho$ for some $\rho\in\mathcal{R}$. That is, the numerators belong to the saturation
@@
With regard to the invariant described by Eq. (168), let us briefly address whether all elements of $H^{2}(G,U(1))$ can actually be achieved by a QCA [recall that this was not necessarily the case for the $H^{1}(G,U(1))$ invariant discussed in $d=0$, for infinite groups $G$]. This invariant characterizes the obstruction in lifting the projective representation $v_{x}$ to a unitary representation. For finite groups, all such invariants can be achieved. Taking the local Hilbert space to be the span of $|g\rangle$ with $g\in G$, and given any 2-coycle $\alpha:G\times G\to U(1)$, we define
$v_{0}(g)=\rho_{\mathrm{reg}}(g)D(g)$ (177)
where $\rho_{\mathrm{reg}}$ is the regular representation and $D(g)$ is the diagonal matrix with elements $\langle h|D(g)|h\rangle=\alpha(g,h)$. Then $v_{0}$ is known to be a projective representation with obstruction $[\alpha]\in H^{2}(G;U(1))$. By a similar construction to the previous paragraph, we can construct a QCA (indeed, a finite depth circuit) from the unitary gate $u_{0,1}$ which conjugate $\rho_{\mathrm{reg}}\otimes\rho_{\mathrm{reg}}$ to $v_{0}\otimes v_{1}$, where $v_{1}$ has obstruction $-[\alpha]$. This gate exists by the vanishing of the obstruction for $v_{0}\otimes v_{1}$. We leave open the precise restrictions on $\omega\in H^{2}(G;U(1))$ imposed by limited ancillae representations or by $G$ being infinite.
The full classifying group is then a subgroup
$\pi_{0}(\mathrm{QCA}_{1}^{\mathcal{R}})\leq H^{2}(G;U(1))\oplus\log\mathcal{R}^{-1}\hat{\mathcal{R}}/X_{G},$ (178)
with only the first summand possibly allowing for this to be a proper subgroup. That is, the projection to the second summand is surjective. We express the second summand with a logarithm just to make the group operation addition. The induced homomorphism from evaluation on a product state, $e_{0}^{*}$, just corresponds to projecting to the first summand, and thus reproduces the group cohomology classification of invertible states in one dimension [66, 93].
@@ E.1 Modified swindle
[where $u_{j,j+1}$ is defined in Eq. (7)] for $t\in[0,\tfrac{1}{3})$,
@@ E.2.1 Invariant
The invariant associated to $v(\theta)$ is a coset [Eq. (141)] of formal ratios of representations with the numerator weighted by the winding numbers. We identify the charge 0 representation with 1 and the charge 1 representation with $x\in{\mathbb{N}}[x,x^{-1}]\cong\mathrm{Rep}(U(1))$. Then the invariant is
@@
Indeed, we believe the odd $W$ phase without symmetry to be the two-tone-driven phase of Refs. [29, 30]. While the solvable limit of the model in Ref. [29] is highly singular, which makes computing invariants from integral formulae inconvenient, the edge phenomenology described in Ref. [29] is consistent with what we expect from having $\tilde{w}_{1}\neq 0\bmod 1$. Namely, the edge dynamics was argued to be anomalous in that it could not be achieved by a zero-dimensional localized [in the sense that a continuous micromotion operator exists as in Eq. (35)] system which did not occur at the edge of an extended one-dimensional system. As the model in Ref. [29] is based around a version of the AKLT model, which can be trivialized when stacked with itself, it is thus reasonable to suppose that $\tilde{w}_{1}=1/2\bmod 1$ for that model, as this is the only element of ${\mathbb{Q}}/{\mathbb{Z}}$ of order 2. Thus, the abstract classification reveals a relationship between the interacting QP pump and Refs. [29, 30] which was not clear previously. It also extends the arguments for stability in Ref. [29] to the nonperturbative regime, where a picture of emergent symmetry (as employed in that reference) becomes inapplicable.